If you put the one card back in before trying it the second time, then the odds are slightly worse; otherwise, the odds are the same - it does not matter if you pull two cards at once or one card twice, subject to the noted limitation.
If you choose two cards from a standard deck what is the probabillity they will be aces?
There are 52 cards in a deck (1-10, plus the Jack, Queen, King, and Ace).
There are four suits in a deck (hearts, spades, clubs, diamonds), and thus, there are 4 possible aces in the deck.
4/52 = 1/13
What is the probability of toxic shock syndrome?
Data on Toxic Shock Syndrome warps probability. As TSS is associated with tampons, which make a lot of money for big companies who sell the most dangerous types of tampons, manufacturers misrepresent TSS statistics - CDC statistics are often quoted but these are grossly inaccurate for a number of reasons. It's estimated that 1 in 700 women will contract TSS at some point in their lives.
What is the probability of the coin landing heads up twice and tails up once when tossed 3 times?
0.666
^^^^
his answer
mine
3 in 8 i believe. which is 0.375
you can get
h,h,h
h,t,h
h,t,t
h,h,t
t,h,h
t,t,h
t,h,t
t,t,t
the ones you are looking for are
h,h,t
t,h,h
h,t,h
What is the probability of a 13 child family having 13 boys?
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
However, if you assume that they are independent events then, given that the probability of a boy is approx 0.52, the probability of 13 boys in a family with 13 children is approx 0.00019.
What is the probability of having three girls and zero boys?
Assuming that having boys and girls are equally likely, then the probability is 1/8.
* * * * *
You also need to assume that the children's genders are independent. They are NOT and depend on the parents' ages and genes.
What is epsilon in statistics?
Epsilon is used with reference to sphericity. It is simply a value by which the degrees of freedom for the test of F ratio if multiplied.
When epsilon is equal to 1 sphericity is met (hence no adjustment). When it is less than 1 the assumption the the covariances and variances are violated. Greenhouse Geisser epsilon is the most common.
When intelligence came into play, life was able to trick itself into believing that causality can be relativised and even overcome. Intelligence uses assumptions to try to comprehent it's environment. In order to make assumptions, a degree of likelyhood has to determined.
Probability is one way of determining likelyhood, as is belief.
well that is not a correct answer
actually in 18th or may be 17th century, some gambling was going in a casino, n there was a game of cards, due to some reason they have to stop the game in middle, the prize amount have to be divided between two parties. the supervisor wasnt sure how to deal with that matter, so he send this problem to his frind who was a famous mathmatician at that time he tried nn tried but failed then he send it to another mathematician who was living somewhere else then he reserch on it, and find a solution... and from that day the probability or statistics born....
What is the probability of becoming male or female?
there are half the chance of a female and half the chance of male so probabilty in both cases is 1/2
What is the similarities between population and sample?
Similarity: Both are counts of people/animals/things.
Difference: Population is the total # of things, while sample is the # of things that you gather data on.
If you pick the right sample size, you can be pretty confident that the results of the sample data is the same as the results of the entire population.
What is standard minute value in garments?
• SMV means Standard Minute Value.
• It is a numerical value which is represented the standard time of a process or operation in a standard environment for standard worker.
• It will be calculated by the following formula (Stop-watch system):
Standard Time = Basic Time + Allowance%
There are many techniques which have been developed to establish standard minute Values, one such technique is Called Pro-SMV , Professional standard minute value, this is a system based on MTM 2 developed by Methods Apparel consultancy India Pvt Limited.
What probability indicates an event that is neither unlikely or likely near?
It is a probability of between 0.3 and 0.7
Assuming the die is six-sided, you would expect to receive 5/6 + 3/6 + 2/6 = 10/6 units payoff each roll.
One would suggest a cost of 2 units per roll to generate profit for the person running the contest.
Will a card drawn from a standard deck be an ace?
There is a 1 in 13 chance
excluding joker cards.
2 to 10 J Q K A = 13
What is the chance of accurate underwater sonar imaging?
Very high.
Not sure you can express it as a probability question though.
The actual definition depends on various factors including frequency (wavelength), scanner geometry and performance, scanning speed and geometry, the image-processing, water depth, scattering and absorption, nature of the objects being scanned, etc., but sonar performance is increasing all the time as the manufacturers all seek to produce better instruments than their rivals!
The z-score must be 1.87: the probability cannot have that value!
What is the probability of getting a sum of 7 or 11 by rolling two dice if the first roll is a 3?
If the first roll is a three, it's impossible to get an eleven (the highest you can get is a 9), so the probability is 0. The way to get a sum of 7 is to add 3 and 4. The probability of getting a 4 on the second die is 1/6. Therefore, by adding the two probabilities of the outcomes (0 and 1/6), you get the answer: 1/6.
How do you express probabilities?
Probabilities are expressed a few different ways, often depending on the means of calculating the probability.
Theoretical probabilities are generally calculated as the number of successes divided by the total number of possibilities. For example, rolling a number greater than 4 on a die has two successes (5 or 6) and six possibilities (1, 2, 3, 4, 5, or 6). Because each of these outcomes is equally likely to occur, the probability is 2/6 = 1/3. Probabilities calculated in this way are typically expressed as fractions, although they can also be expressed as decimals or percents (more commonly referred to as the chance of an event occurring).
Probabilities based on the normal curve, models, random variables and the like are generally expressed as decimals, although percentages are also acceptable.
Note that for all of these forms, probabilities are always between 0 and 1 inclusive (0% and 100%). A probability of 0 (0%) has absolutely no chance of occurring while a probability of 1 (100%) means an event is certain to occur. Probabilities in between represent a certain degree of likelihood of the event occurring.
What is the probability of rolling a four on one roll of a dice?
It is going to be 1 over 6 because there is only 6 numbers on a dice and only one 5.